triangle EBT is similar to SDA. if the perimeter of EBT is 63, the perimeter of SDA is 21. BT=6x, and DA=x+3, find BT and DA
Answer by math_tutor2020(3817)


The linear scale factor from EBT to SDA is 1/3 because 21/63 = 1/3.
The side lengths of SDA are 1/3 as long compared to the side lengths of EBT.

DA/BT = 1/3
(x+3)/(6x) = 1/3
3(x+3) = 6x*1
3x+9 = 6x
9 = 6x-3x
9 = 3x
x = 9/3
x = 3

BT = x+3 = 3+3 = 6
DA = 6x = 6*3 = 18